When a Fr-223 atom with 87 protons undergoes TWO successive alpha decays, what is the daughter nucleus after the second decay?

Rn-222 with 86 protons

Po-219 with 84 protons

Th-227 with 92 protons

Po-215 with 84 protons

Bi-215 with 83 protons
Answer #1

d- particle 2 prown +2 numn -t fr-223 (328) to Merala P= 87 -2=85 A 2223-42219 - Mereela p=85-2=83 A = 219-4=215 So, our rohn